# Who Does the Giant Number Pile Like Best: Analyzing Fairness in Hiring Contexts

### v1

@@ 3.3 Non-Demographic Perturbations
We do two non-demographic perturbation experiments for retrieval to assess baseline sensitivity of embedding models that is not due to demographics.
Within Group Name Perturbations
As a baseline comparison to performing name perturbations between different demographic groups (e.g., White female $\rightarrow$ Black Female), we assess whether models are sensitive to same group demographic perturbations (e.g., White female $\rightarrow$ White Female). By doing so, we disentangle how much bias is due demographic changes vs. overall model sensitivity to name changes.
To control for the effects of frequency (Ethayarajh et al., 2019), we bin names in each demographic group according to their frequency in the Pile dataset (Gao et al., 2020), and match names based on the bin. Therefore, name perturbations should result in a similarly occurring name within the same demographic group.
@@
We also assess whether model rankings are sensitive to non-name perturbations. This allows us to examine whether models lack robustness more broadly, or mainly exhibit brittleness with respect to name changes.
We test two perturbation types: (1) random character swapping, which does not impact readability or comprehension in a meaningful way and (2) replacing new lines in the resume with a single space instead, which targets formatting without modifying content. For character swapping, we choose 10 random characters in the resume and swap based on neighboring keys to simulate typos.
@@ 3.4 Models
We generate summaries using a combination of closed and open state-of-the-art language models: GPT-4o (OpenAI, 2024), Command-R (Cohere, 2024), and Mixtral 8x7B (Jiang et al., 2024). For summary instructions, we vary the generation length (100 and 200 words) as well as the point of view (first and third person) specified in the prompt. We generate summaries with temperatures of 0.0 and 0.3.
@@
For retrieval, we select four popular dense embedding models used in retrieval augmented generation (RAG) systems (Lewis et al., 2020): OpenAI’s text-embedding-3-small and text-embedding-3-large models, Cohere’s embed-english-v3.0, and Mistral’s mistral-embed.
(a) Top-5%
(b) Top-10%
Figure 2: Non-uniformity metric for top-5 and top-10% of retrieved resumes. Separated (sep) measures the % of individual job posts where the top-$x$of resumes are non-uniformly distributed, while pooled (pool) measures the % of occupations where the top-$x$of resumes across job posts for that occupation are non-uniformly distributed.
@@ 4.1 Summarization
We analyze whether generated summaries differ meaningfully across demographic groups (RQ1) by examining invariance violations, i.e., the percentage of t-tests that yield significant differences in our automated measures. We measure across summaries with different characteristics (length, point of view, and temperature). Figure 1 displays results for generated resumes, grouped by completion model and demographic perturbation.
Mixtral 8x7B exhibits the highest average invariance violations (8.59%), followed by GPT-4o (3.91%) and Command-R (2.74%). All models violate invariance much more for resumes that differ by race vs. gender. For example, the average invariance violations for race perturbations is 15.62% for Mixtral 8x7B, as compared to only 1.56% for gender. In summary, we observe that models exhibit some but not considerable discrepancies between generated summaries for different demographic groups, with different models having potentially large differences in sensitivity.
@@ 4.2 Retrieval
We now shift to a retrieval setting and ask: do models exhibit fairness issues in resume retrieval? Our analysis tackles this question from both a distributional (RQ2) and robustness (RQ3) perspective.
@@ 4.2.1 Non-Uniformity
We compute non-uniformity (i.e., how often top retrieved resumes have non-uniform demographic distributions). Non-uniformity differs considerably across models, choice of top-$x$ percent, and whether or not resumes are pooled across occupations (Figure 2).
@@
We observe that embed-english-v3.0 exhibits the highest non-uniformity on average, where 6.90% of job posts and 45.45% of occupations have non-uniformly distributed resumes. We see that increasing top-$x$ from 5% to 10% and pooling resumes both yield higher non-uniformity across all models. In particular, pooling resumes across occupations can produce massive changes – for example, we see that on average across models, non-uniformity goes from 3.66% $\rightarrow$ 30.68% after pooling, resulting in a nearly 750% increase. However, this reflects a sensitivity in the metric itself more than a change in the shape of the distribution. Increasing top-$x$ and pooling both increase sample size, which can lead to rejecting the null hypothesis that the distribution is uniform in cases where the null hypothesis previously failed to be rejected.
Different models demonstrate distinct patterns of bias. For example, when considering the highest represented groups in the top-10% of resumes with respect to each job post, White females are the top group in 37 job posts (48.05%), while Black males are the top group in only 3 job posts (3.90%) for embed-english-v3.0. In contrast, for mistral-embed, White males are the top group in 56 job posts (72.73%), while White females are the top group in only 4 job posts (5.19%). We cannot determine why such differences exist without access to datasets or training details, but it is notable and somewhat surprising that models do not uniformly prefer one demographic group over the other.
@@ 4.2.2 Exclusion
We compute exclusion (i.e., how often top retrieved resumes are excluded from the set of top-$n$ resumes after demographic perturbations). We find that all models display notable sensitivity to gender and race name perturbations, as shown in Figure 3. When considering top-5 resumes, we find that models tend to exclude perturbed resumes nearly half the time (45.75% on average).
Across both gender and race name perturbations, and different $n$ values, text-embedding-3-small and text-embedding-3-large have the highest exclusion, while embed-english-v3.0 consistently has the lowest exclusion. As expected, exclusion lowers as $n$ increases, since larger $n$ values are less restrictive and consider a larger set of retrieved resumes. That being said, exclusion for $n=100$ is still considerable, as all models have exclusion > 12%. [^6] In practice we expect $n$ to be low for filtering candidates.
For exclusion, models have similar sensitivity to both gender and race perturbations. Overall, average exclusion for gender is 31.78% on generated resumes (25.47% on Kaggle resumes) vs. 31.66% on generated resumes (25.40% on Kaggle resumes) for race. [^7] Kaggle resumes exhibit similar patterns but lower exclusion magnitude than generated resumes. This is likely because generated resumes are tech-focused and closer in content. These results contrast our summarization findings, where models show greater invariance violations on race vs. gender perturbations. Our analysis reveals that the set of top retrieved resumes with respect to a job posting is highly brittle, as merely altering the demographic on resumes (through names only) often results in otherwise identical resumes dropping out of the top-$n$ results.
@@
We can also partition the results based on direction of perturbation (Figure 4). We find that the magnitude of difference exceeds a threshold of 5% (e.g., the absolute difference in exclusion for M $\rightarrow$ F vs. F $\rightarrow$ M or W $\rightarrow$ B vs. B $\rightarrow$ W for the same model, $n$, and resume type is $>5$) in 26/48 cases across generated and Kaggle resumes, indicating models often exhibit higher sensitivity to one direction of perturbation over the other. In particular, the gender directional difference is notable for mistral-embed, going from 63.28% for M $\rightarrow$ F to 27.93% for F $\rightarrow$ M, for generated resumes with $n=5$ (Figure 4).
@@
Figure 5: Deltas (differences) in exclusion metric for retrieval after performing demographic perturbations with names as well as extracurricular information as compared to only names, separated by direction. As expected, adding extracurricular information increases sensitivity to perturbations.
We also observe that models exhibit opposite directional trends for gender and race in Figure 4. For gender, all models except mistral-embed are more sensitive when perturbing female names (marginalized) to male names (non-marginalized); for race, all models except mistral-embed are more sensitive when perturbing White names (non-marginalized) to Black names (marginalized). These results highlight an asymmetry in how models handle different types of demographic changes.
@@
Figure 5 shows that, as expected, models tend to be more sensitive when perturbing extracurricular information (e.g., awards, clubs, volunteering, etc.) in addition to names. On average, across models and $n$ values, we observe the following increases in exclusion: 9.35% for M $\rightarrow$ F, 8.06% for F $\rightarrow$ M, 16.41% for W $\rightarrow$ B, and 2.90% for B $\rightarrow$ W.
For gender, adding extracurricular information results in comparable increases in exclusion for both directions. In contrast, adding extracurricular information for race results in highly asymmetric increases in exclusion. W $\rightarrow$ B averages more than 5x the increase of B $\rightarrow$ W changes. We observe that adding extracurricular information results in non-uniform increases to exclusion, which suggests that models may encode and utilize various types of demographic signal differently. This finding is notable given that prior work often examines a single way of encoding demographics, overlooking how various signals interact and compound.
@@
We consider two sets of perturbation analyses that are non-demographic: (1) How sensitive are models to within-group name perturbations? and (2) How sensitive are models to non-name perturbations? Even when perturbing names within the same demographic group, models surprisingly exhibit similar levels of sensitivity to those observed with demographic name perturbations (Figure 6(a)).
(a) Within Group Name Perturbation
(b) Non-name Perturbation
Figure 6: Exclusion metric for retrieval after performing non-demographic perturbations (i.e., within group name changes - left, and modifying spacing and adding typos - right). Results shown for $n=5$.
For non-name perturbations, we find that models are extremely sensitive to both spacing and typos, although to a lesser extent than name perturbations. As shown in Figure 6(b), most models demonstrate higher sensitivity to spacing than typos, though there is surprising sensitivity to both. In particular, mistral-embed excludes resumes from the top-5 set 72.76% of the time solely based on spacing, which indicates that simple formatting choices can have a massive impact on fairness (in this case, much more than names). Generated resumes display nearly twice the sensitivity to spacing changes compared to web-scraped Kaggle resumes (27.98% vs. 15.35% averaged across models and $n$ values), likely due to their more structured and formatted nature. In summary, we observe that retrieval models lack robustness more generally, which might contribute to fairness-related issues.
@@ 5 Discussion
Our results highlight important issues about model fairness and brittleness in real-world contexts. First, we observe that model rankings in a retrieval setting are impacted considerably by both subtle demographic and non-demographic changes. In practice, these differences would lead to unintended exclusion, with candidates being eliminated from consideration at an early stage. We also observe that subtle demographic differences in resumes can alter the way candidates are discussed in generated summaries. As a result, even candidates who make it past the initial resume screening stage may be portrayed differently based on demographic attributes. These differences can impact downstream decision-making, since recruiters might rely solely on summaries to shortlist candidates.
Additionally, it is important to consider compositional effects when combining components (i.e., ranking $\rightarrow$ summarization), since biases can compound due to the sequential nature of these tasks. Therefore, the candidate pool may (1) leave out qualified candidates through retrieval bias, as well as (2) differentially represent candidates through summarization bias. Applying this to our results, at worst case, we find that modifying racial indicators on resumes from W $\rightarrow$ B (using both names and extracurricular information) can result in roughly (1) 70% of candidates being filtered out at the resume screening stage, and (2) 15% of remaining candidates being depicted in less preferable ways.
Our analysis also reveals that all models are sensitive to non-demographic perturbations. These findings suggest that model unfairness may partially stem from more general robustness issues, rather than solely from encoded biases. While these insights do not change the real-world impacts of discriminatory outcomes, understanding that disparate treatment can arise from multiple factors—including but not limited to undesirable biases—provides a more complete and nuanced picture for addressing fairness issues. Moreover, given that retrieval models are commonly used in RAG systems, these issues may extend to various settings and applications beyond HR.
@@ 6 Conclusion
We examine fairness in LLM-based hiring systems by analyzing two key components: applicant ranking and summary generation. We propose two new metrics to measure fairness in retrieval and generative settings, and find that a hiring pipeline with these components produces biased outcomes, particularly during the initial retrieval phase. In addition, we see unexpected sensitivity to minor non-demographic changes, indicating a lack of overall robustness that may contribute to unfair outcomes. We hope that our work underscores the need for more real-world evaluations of LLM harms, especially in other high-stakes domains.
@@
Our analysis focuses exclusively on English resumes and job posts, which poses limitations. Future research should investigate fairness considerations in multilingual settings and examine whether our conclusions hold across various languages. Additionally, cultural norms likely influence how candidates present themselves and describe their professional experience, qualifications, and achievements. Understanding these nuances is crucial for evaluating and developing hiring systems that serve diverse global talent pools.
While our analysis examines whether hiring systems behave differently for various gender (male and female) and racial (White and Black) groups, it is meant to be illustrative rather than exhaustive and only covers a subset of gender and racial identities. We only consider binary gender biases, and exclude non-binary gender biases from our analysis (e.g., if a candidate explicitly declares pronouns on their resume). We also only focus on Black and White racial groups, since this is a common emphasis in fairness studies. In addition, future research should expand beyond these commonly investigated biases and analyze to what extent other types of demographic information (e.g., age and nationality) impact LLM fairness in hiring.
@@ Appendix A Appendix
A.1 Names

### v2

@@ 3.3 Non-Demographic Perturbations
We conduct two non-demographic perturbation experiments for retrieval to assess the baseline sensitivity of embedding models that is not due to demographics.
Within-Group Name Perturbations
As a baseline comparison to performing name perturbations between different demographic groups (e.g., White female $\rightarrow$ Black Female), we assess whether models are sensitive to within-group demographic perturbations (e.g., White female $\rightarrow$ White Female). By doing so, we disentangle how much bias is due to demographics vs. model sensitivity to name changes. To control for the effects of frequency (Ethayarajh et al., 2019), we bin names in each demographic group according to their frequency in the Pile dataset (Gao et al., 2020), and match names based on the bin. [^7] We use the What’s in My Big Data tool (Elazar et al., 2023) to obtain frequencies.
@@
We assess whether model rankings are sensitive to non-name perturbations. This allows us to examine whether models lack robustness more broadly. We test two perturbation types: (1) random character swapping, which does not impact readability or comprehension in a meaningful way [^8] We choose 10 random characters in the resume and swap with neighboring keys to simulate typos. and (2) replacing new lines in the resume with a single space instead, which targets formatting without modifying content.
(a) Gender
(b) Race
Figure 2: Summarization Results: Invariance violations for generated summaries, separated by completion model and perturbation type. Results are shown across 5 runs. Left 3 models are considered "smaller" models, right 3 models are considered "larger" models.
@@ 3.4 Models
We generate summaries using closed and open state-of-the-art LLMs: GPT-4o (OpenAI, 2024), Command-R (Cohere, 2024), Mixtral 8x7B and Mistral Large (Jiang et al., 2024), and Llama 3.1 8B and Llama 3.3 70B (Meta AI, 2024). For summary instructions, we vary the generation length (100, 200 words) as well as the point of view (first, third person) specified in the prompt. We generate summaries with temperatures of 0.0 and 0.3. To account for stochasticity in generations, we generate each summary five times.
@@
For retrieval, we select four popular dense embedding models used in retrieval augmented generation (RAG) systems (Lewis et al., 2020): OpenAI’s text-embedding-3-small and text-embedding-3-large, Cohere’s embed-english-v3.0, and Mistral’s mistral-embed.
@@ 4.1 Summarization
We analyze whether generated summaries differ meaningfully when applying gender and race perturbations (RQ1) by examining invariance violations, i.e., the percentage of t-tests that yield significant differences in our automated measures. We measure violations separately for summaries with different characteristics (length, point of view, and temperature). Figure 2 displays results grouped by completion model and perturbation type.
All models violate invariance much more for resumes that differ by race as opposed to gender. In fact, gender invariance violations are zero or near zero for all models. In contrast, all models except Command-R exhibit invariance violations with respect to race, with Mixtral 8x7B exhibiting violations 16.76% of the time on average. Our results also provide some indication that smaller models are more susceptible to violations. In summary, we observe that models exhibit some but not considerable discrepancies between generated summaries for different demographic groups, with minimal differences for gender perturbations.
@@ 4.2 Retrieval
Moving to retrieval, we ask: Do models exhibit fairness issues in selecting resumes? Our analysis tackles this question from distributional (RQ2) and robustness (RQ3) perspectives.
@@ 4.2.1 Non-Uniformity
To answer this question, we compute non-uniformity (i.e., how often top retrieved resumes have non-uniform demographic distributions). All models disparately retrieve resumes across demographic groups, consistent with the findings of Wilson and Caliskan (2024). That being said, non-uniformity differs considerably across models, choice of top-$x$ percent, and pooling of resumes across occupations (Figure 3).
We observe that embed-english-v3.0 exhibits the highest non-uniformity on average, with 6.90% of job posts and 45.45% of occupations having non-uniformly distributed resumes. Increasing top-$x$ from 5% to 10% and pooling resumes both yield higher non-uniformity across all models. In particular, pooling resumes by occupation can produce massive changes; on average across models, non-uniformity goes from 3.66% $\rightarrow$ 30.68%. This reflects sensitivity in the metric itself more than a change in the shape of the distribution. [^9] Increasing top-$x$ and pooling both increase sample size, which can lead to rejecting the null hypothesis in cases where the null hypothesis previously failed to be rejected.
(c) Top-5%
(d) Top-10%
Figure 3: Non-uniformity metric for top-5 and top-10% of retrieved resumes. Separated (sep) measures non-uniformity at a job post level, while pooled (pool) measures it at an occupation level by pooling results across job posts for a given occupation.
@@
Different models show distinct patterns of bias: the non-uniformity privileges different demographic groups. For example, in the top-10% of resumes from embed-english-v3.0, White females are the top group 48.05% of the time, compared to 3.90% for Black males. In contrast, for mistral-embed, White males are the top group 72.73% of the time, compared to 5.19% for White females. Reasons for these differences are unclear without access to dataset and training details but it is notable and surprising that models do not consistently favor the same demographic group.
@@ 4.2.2 Exclusion
We compute our proposed metric, exclusion (i.e., how often top retrieved resumes are excluded from the set of top-$n$ resumes after perturbations), and find that all models display notable sensitivity to gender and race name perturbations (Figure 4). When considering the top-5 resumes, we find that models tend to exclude perturbed resumes nearly half the time (45.75% on average).
Across both gender and race name perturbations, and different $n$ values, text-embedding-3-small and text-embedding-3-large have the highest exclusion, while embed-english-v3.0 consistently has the lowest exclusion. As expected, exclusion lowers as $n$ increases, since larger $n$ values are less restrictive and consider a larger set of retrieved resumes. That being said, exclusion for $n=100$ is still considerable, as all models have exclusion >> 12%. [^10] In practice we expect $n$ to be low for filtering candidates.
@@
In contrast to our summarization findings, where models show greater invariance violations for race vs. gender perturbations, models have similar sensitivity to gender vs. race perturbations for exclusion. Overall, average exclusion for gender is 31.78% on generated resumes (25.47% on Kaggle resumes) vs. 31.66% on generated resumes (25.40% on Kaggle resumes) for race. [^11] Kaggle resumes exhibit similar patterns to generated resumes, but are lower in exclusion magnitude. This is likely because generated resumes are tech-focused and more overlapping in content. Our analysis reveals that the set of top retrieved resumes with respect to a job posting is highly brittle, as merely altering the demographic with names often results in otherwise identical resumes dropping out of the top-$n$ results.
@@
Figure 6: Deltas (differences) in exclusion metric for retrieval after performing demographic perturbations with names + extracurricular information vs. names only. As expected, adding extracurricular information increases sensitivity to perturbations.
We partition the results based on the perturbation direction (Figure 5), and find that models often exhibit higher sensitivity to one direction of perturbation over the other. In particular, the gender directional difference is notable for mistral-embed, going from 63.28% for M $\rightarrow$ F to 27.93% for F $\rightarrow$ M, for generated resumes with $n=5$. We also observe that models exhibit opposite directional trends for gender and race. For gender, all models except mistral-embed are more sensitive when perturbing female names (marginalized) to male names (non-marginalized). On the other hand, for race, all models except mistral-embed are more sensitive when perturbing White names (non-marginalized) to Black names (marginalized). These results highlight an asymmetry in how models handle various demographic changes.
@@
Figure 6 shows that models tend to be more sensitive when perturbing extracurricular information in addition to names. On average, we observe the following increases in exclusion: 9.35% for M $\rightarrow$ F, 8.06% for F $\rightarrow$ M, 16.41% for W $\rightarrow$ B, and 2.90% for B $\rightarrow$ W.
For gender, adding extracurricular information results in comparable increases in exclusion for both directions. In contrast, adding extracurricular information for race results in highly asymmetric increases. W $\rightarrow$ B averages more than 5x the increase of B $\rightarrow$ W changes. We observe that adding extracurricular information results in non-uniform increases to exclusion, which suggests that models may encode and utilize various types of demographic signal differently. This finding is notable given that prior work often examines a single way of encoding demographics, overlooking how various signals interact and compound.
(a) Within-Group Name, $n=5$
(b) Non-name, $n=5$
Figure 7: Exclusion metric for retrieval after performing non-demographic perturbations (i.e., within group name changes - left, and modifying spacing and adding typos - right).
@@
To disentangle fairness from robustness issues, we consider two sets of perturbation analyses that are non-demographic: (1) How sensitive are models to within-group name perturbations? and (2) How sensitive are models to non-name perturbations? Even when perturbing names within the same demographic group, models surprisingly exhibit highly similar levels of sensitivity to those observed with gender and race name perturbations (Figure 7(a)).
We find that models are extremely sensitive to both spacing and typos, but to a lesser extent than names. As shown in Figure 7(b), most models demonstrate higher sensitivity to spacing than typos, though there is surprising sensitivity to both. In particular, mistral-embed excludes resumes from the top-5 set 72.76% of the time solely based on spacing, which indicates that formatting can have a massive impact on fairness (in this case, much more than names). Generated resumes display nearly twice the sensitivity to spacing changes compared to web-scraped Kaggle resumes (27.98% vs. 15.35% averaged across models and $n$ values), likely due to their structured formatting. In summary, we observe that retrieval models lack overall robustness, which has fairness implications.
@@ 5 Discussion
Our results highlight failures of both fairness and robustness in LLMs in hiring contexts. We differ from prior work on LLM fairness in summarization (Zhang et al., 2024) and ranking (Xu et al., 2024) in that our evaluations are grounded in real-world applications, and this reveals novel insights that have ecological validity. First, we observe that model rankings in a retrieval setting are impacted considerably by both subtle demographic and non-demographic changes. In practice, these differences would lead to unintended exclusion, with candidates being eliminated from consideration during initial screening. We also observe that subtle demographic differences in resumes can alter the way candidates are discussed in generated summaries. As a result, candidates who make it past the initial resume screening stage may be portrayed differently based on demographic attributes, which can impact downstream decision-making.
Additionally, it is important to consider compositional effects when combining components, since biases can compound due to the sequential nature of these tasks. The resulting candidate pool may (1) leave out qualified candidates through retrieval bias, and (2) differentially represent candidates through summarization bias. Applying this to our results, in the worst case, we find that modifying racial indicators on resumes (using names + extracurricular information) can result in roughly (1) 70% of candidates being filtered out at the resume screening stage, and (2) 20% of remaining candidates being depicted in less preferable ways.
Our analysis also reveals that all models are sensitive to non-demographic perturbations, suggesting that model unfairness may partially stem from more general robustness issues, rather than encoded biases alone. These perturbations still result in disparate outcomes, but with different underlying causes. While these insights do not change discriminatory impact, understanding that disparate treatment can arise from small input changes, both demographic and otherwise, provides a more complete picture for addressing fairness issues.
@@ 6 Conclusion
We examine allocational fairness in LLM-based hiring systems by analyzing two key components: applicant ranking and summary generation. To support systematic measurement and mitigation of fairness issues, we release a benchmark dataset and introduce an evaluation framework with new metrics. [^12] The data and code can be found at: https://github.com/preethisesh/hiring_fairness. We find that a hiring pipeline consisting of these two stages produces biased outcomes, particularly during the retrieval phase. In addition, models show unexpected sensitivity to minor non-demographic changes, revealing a lack of overall robustness that may contribute to unfair outcomes. These findings underscore the need for targeted strategies to improve the fairness of LLM-based hiring, and the importance of realistic, application-grounded evaluations of LLM harms.
@@
Our analysis focuses exclusively on English resumes and job posts. Future research should investigate fairness considerations in multilingual settings and examine whether our conclusions hold across various languages. Additionally, cultural norms likely influence how candidates present themselves and describe their professional experience, qualifications, and achievements. Understanding these nuances is crucial for evaluating and developing hiring systems that serve diverse global talent pools. Since we are releasing our code and datasets, researchers in other regions will be able to expand our work as well.
While our analysis examines whether hiring systems behave differently for various gender (male and female) and racial (White and Black) groups, it is meant to be illustrative rather than exhaustive and only covers a subset of gender and racial identities. We only consider binary gender biases, and exclude non-binary gender biases from our analysis, since this information cannot be inferred from a name. While candidates may explicitly declare pronouns on resumes, we do not observe this in the resumes we collect, so we do not vary them. In addition, we only focus on Black and White racial groups, since this is a common emphasis in fairness studies, and only to do so in the context of US names. We hope future work expands beyond these commonly investigated biases and analyzes the extent to which other types of demographic information (e.g., age and nationality) impact LLM fairness in hiring.
Moreover, although the way we handle name perturbations is standard practice in NLP fairness literature, we acknowledge that names can encode demographic axes beyond gender and race, including age, class, and region. These signals are more subtle and challenging to isolate, making it difficult in practice to vary only a single dimension at a time. It is worth noting that we control for other factors such as name frequency to reduce potential confounds.
@@ Appendix A Appendix
A.1 Additional Background and Related Work
For background on the allocational fairness of LLMs in high-stakes domains, please see the Introduction.
Name Perturbations
Performing name perturbations to study fairness is common practice in NLP fairness literature (Webster et al., 2021; An and Rudinger, 2023; Steen and Markert, 2024; Wan et al., 2023; An et al., 2024). We go beyond this by perturbing resumes with extracurricular information, as done in Glazko et al. (2024), but largely focus on names because it is common practice. It is worth pointing out that Gautam et al. (2024) highlight limitations around inferring sociodemographic groups from names, such as poor validity. We try to account for some of these concerns by using the carefully curated names from Yin et al. (2024).
Fairness Definitions
We draw connections between the metrics we use and traditional ML fairness metrics (Mehrabi et al., 2021). Non-uniformity is connected to statistical parity, which is satisfied if the probability of a prediction is independent of demographic group. We adapt this idea by evaluating for non-uniformity in the demographic distribution of top-x%. Exclusion bears resemblance to both individual fairness (Dwork et al., 2012), which assesses whether similar individuals are treated similarly, and counterfactual fairness (Kusner et al., 2017), which assesses whether outcomes are consistent for counterfactual individuals. Similarly, exclusion measures the stability of rankings under demographic perturbations.
Fairness in Summarization and Ranking
Several studies have identified biases in LLM-generated summaries (Shandilya et al., 2018; Guo et al., 2023; Zhang et al., 2024; Li et al., 2025), but they do not conduct application-grounded evaluations or consider allocational harms. A few recent works have also studied the fairness of LLMs in ranking (Wang et al., 2024; Xu et al., 2024). Similarly, these works mainly focus on traditional retrieval tasks such as article relevance, rather than real-world LLM usage in high-stakes domains like hiring.
A.2 Focus on Evaluation over Mitigation
Given the lack of work on investigating allocational harms in LLM-based hiring systems, our main goal is to establish a comprehensive benchmark of fairness risks. Benchmarking is necessary to first understand fairness issues, and mitigation is the natural next step. Meaningful progress towards mitigation cannot be made without proper evaluation and metrics—a shared framework is necessary to compare the performance of mitigation methods and track improvements. We will make our data and code available to the community, which will enable testing various mitigation approaches.
A.3 Names